# Coarse-graining nonequilibrium diffusions with Markov chains

### v1

@@ 4.1.1. The Ornstein-Uhlenbeck process
The Ornstein-Uhlenbeck (OU) process is a multivariate diffusion process that models interacting degrees of freedom under both friction and fluctuations [Godreche2018OU]. It is given by the SDE,
@@
where $B\in\mathbb{R}^{d}\times\mathbb{R}^{d}$ is the friction matrix and $D=\frac{1}{2}\Sigma\Sigma^{\top}$ is the diffusion matrix. If all eigenvalues of $B$ have positive real-parts, the process converges to a zero-mean Gaussian stationary density given by [^8] As we denote the stationary density as $\pi$, we use $\bm{\pi}$ to denote the mathematical constant.,
@@
which corresponds to the multivariate normal distribution $\mathcal{N}_{d}(0,S)$, where $S$ is the covariance matrix satisfying the Lyapunov equation,
@@
In general, steady-states of the OU are out of equilibrium. More specifically, the steady-state is in equilibrium if and only if $BD=DB^{\top}$ which implies $S=B^{-1}D$ [lax1960fluctuationsnonequilibrium]. Otherwise, the stationary probability flux is given by,
@@
where $\mu=DS^{-1}-B$, which then allows for an explicit calculation of the EPR. We first define Q=BSDQ=BS-D, which implies that $\mu=-QS^{-1}$ and $Q=-Q^{\top}$ from the Lyapunov equation, Eq. (42). Using the definition of the EPR from Eq. (9), we have that,
(44) $\Phi$ $=\langle x^{\top}\mu^{\top}D^{-1}\mu x\rangle,$
where $\langle\cdot\rangle$ represents expectation. We can then use that $\langle x^{\top}Ax\rangle=\operatorname{Tr}(SA)$ for Gaussian $x$ with covariance $S$ to obtain,
(45) $\Phi$ $=\operatorname{Tr}(S\mu^{\top}D^{-1}\mu)$
$=-\operatorname{Tr}(S^{-1}QD^{-1}Q)$
$=\operatorname{Tr}(B^{\top}D^{-1}Q)-\underbrace{\operatorname{Tr}(S^{-1}Q)}_{=0}$
$=\operatorname{Tr}(B^{\top}D^{-1}Q),$
where $\operatorname{Tr}(S^{-1}Q)=0$ as the trace of a symmetric matrix times an antisymmetric matrix is zero [DaCosta_2023].
@@ 4.1.2. The stochastic Hopf oscillator
The Hopf oscillator [^9] Also known as the Stuart-Landau oscillator [Matthews1990phasediagram]. (HO) is a nonlinear process given by the equations,
@@
where $\omega$ is the natural frequency and $\sigma$ is the noise intensity. In the absence of noise, for $a<0$ the system decays to the origin which is an attracting fixed point. At $a=0$, the system goes through a Hopf bifurcation, resulting in limit-cycle oscillations for $a>0$ [strogatz2008nonlinear]. This system has become a useful model of neural dynamics [nartallokalu2025review], and hair-cell bundles [Sheth2018hair].
To study the NESS of this process in the presence of additive noise, we consider the FP equation in an alternative form,
@@
In order to solve for the steady-state, we use that the limit-cycle behaviour of the deterministic system is radially symmetric, and thus assume that the stationary density is also i.e. $\frac{\partial p}{\partial\theta}=\frac{\partial^{2}p}{\partial\theta}=0$, which implies the distribution takes the form,
@@
which is solved by the Boltzmann-Gibbs density [^10] It is not immediately obvious why a solution of the form of Eq. (52) solves Eq. (51). However, it is intuitive that the irreversible, oscillatory circulation is driven by the $\theta$ dynamics, which are absent here. This suggests that the dynamics in $r$ may be of Smoluchowski form. We can see that $f^{r}=-\frac{dU}{dr}$ for a potential $U=\frac{1}{2}r^{2}(a-\frac{1}{2}r^{2})$. All that remains is to check that $\pi(r)=\exp\left(-2U/\sigma^{2}\right)$ solves Eq. (51). Whilst the stationary density is a Boltzmann-Gibbs density, the steady-state is in nonequilibrium.,
@@
where $Z$ is the normalising factor [^11] An alternative approach to solving for the stationary density can be found for the case $a=\omega=1$ in Ref. [Yuan2017decomposition].,
@@
Given the stationary density, we can perform the HHD to obtain,
@@ 4.1.3. Numerical experiments
Next, by varying the parameter $\theta$, we can vary the EPR, $\Phi=\theta^{2}$, of the underlying OU process. Fig. 6 shows the EPR of the coarse-grained process for a range of $\Delta x$ values as a function of $\theta$. Whilst all the coarse-grained dynamics follow the correct trajectory, increasing the EPR with $\theta$, the smaller step-sizes more closely approximate the true EPR. Moreover, we notice that the EPR of the coarse-grained dynamics may exceed the true EPR. Unlike the EPR of a diffusion observed at a coarse-grained level, the Markov process derived here is not subject to the inequality proved by Esposito in Ref. [Esposito2012coarsegraining], where it would necessarily lower-bound the true EPR.
@@ 4.2.1. The stochastic van der Pol oscillator
The van der Pol oscillator (VDP) is a nonlinear, non-conservative oscillator that converges to an asymmetric limit-cycle in the absence of noise (see Fig. 9) [strogatz2008nonlinear]. We consider a modified, stochastic VDP of the form,
@@
For models without an exact stationary density, it is worth additionally verifying that this is the stationary behaviour seen in the long-time-limit, by numerically integrating the FP equation directly. Using the Crank-Nicholson scheme,
@@
where $B=I-\frac{\Delta t}{2}L$ and $A=I+\frac{\Delta t}{2}L$ for a time-step $\Delta t$ [LeVeque2007finitedifference]. We begin with a initial Gaussian density with zero mean and isotropic deviation $\sigma=1$ and integrate up to $T=1.4$ with $\Delta t=0.2$ and $\Delta x=0.05$. As shown in Fig. 10, we see that the FP solution converges to the stationary distribution calculated directly.
@@ 4.2.2. A pair of frustrated Kuramoto oscillators
The Kuramoto oscillator is a model that described the phase of interacting limit-cycle oscillators [Kuramoto1984oscillations], and is typically used to model synchronisation. It has found a range of applications describing neural dynamics [nartallokalu2025review], the flashing of fireflies, or the voltage oscillations of Josephson junctions [strogatz2008nonlinear].
We first consider the case of a single, isolated Kuramoto oscillator,
@@
oscillating at a natural frequency of $\omega>0$ where $\theta(t)\in[0,2\bm{\pi})$. Unlike the other systems we considered, the Kuramoto model is not defined on $\mathbb{R}^{n}$, and instead has a periodic boundary. As a result, even in a single dimension, it can be in a NESS as there can be non-zero flux at the boundary [^14] The result in Sec. 2 refers to 1D diffusions on $\mathbb{R}$, where the flux must decay at the boundary to guarantee stationarity.. More explicitly, we arrive at the FP equation,
@@
Typically, Kuramoto oscillators are placed on a network defined by an adjacency matrix $A=(A_{ij})$ with $A_{ij}$ representing the connection strength from oscillator $j$ to $i$. In the absence of internal oscillations, the dynamics are given by,
(61) $d\theta_{i}(t)$ $=\sum_{j}A_{ij}\sin(\theta_{j}-\theta_{i})\;dt+\sigma\;dW_{i}(t).$
When the connections are symmetric, $A=A^{\top}$, then this is a Smoluchowski dynamic, $d\theta_{i}=-\partial_{\theta_{i}}V\;dt+\sigma\;dW_{i}(t)$, for the potential,
(62) $V=-\frac{1}{2}\sum_{i,j}A_{ij}\cos(\theta_{j}-\theta_{i}).$
As a result, it converges to an ESS given by a Boltzmann distribution, $\pi\propto\exp(-2V/\sigma^{2})$. Internal oscillations and asymmetric interactions cause the process to converge to a NESS without a closed form distribution [fruchart2021nonreciprocal].
Here, we focus on a system with a pair of Kuramoto oscillators with additive noise. In addition, their interaction is frustrated, another feature that can lead to nonequilibrium behaviour [Gupta2014kuramoto]. We consider,
@@
In the first panel of Fig. 14, we calculate the EPR as a function of various values of the coupling strengths. Whilst the system remains out of equilibrium for symmetric coupling, due to the internal oscillations, asymmetry in coupling strength drives the system further from equilibrium, in keeping with previous results on non-reciprocity in collective dynamics [fruchart2021nonreciprocal, nartallokaluarachchi2024broken]. On the other hand, from the second panel, we see that only the absolute value, not the asymmetry, of natural frequencies drives the system from equilibrium, by increasing the speed of the collective oscillation. Finally, in the third panel, we investigate the effect of frustration and noise. Unsurprisingly, noise reduces the EPR by increasing reversible motion. Interestingly, the effect of frustration does not appear to be constant with respect to noise, with the EPR-maximising value of $\alpha$ varying slightly with the noise intensity, $\sigma$.
@@ 5. Statistical inference for discrete-state processes
When analysing NESS in physical and biological systems, we typically do not have access to the function forms of the drift or diffusion. Instead, we must analysis a diffusive process from observations of stochastic trajectories. In order to analyse the observed trajectories of a diffusion, it is common to infer a stochastic model. Whilst some methods attempt to directly infer a SDE [Frishman2020stochasticforce, Friedrich2011complexity, bruckner2020inferring, elbeheiry2015inferencemap], we will instead focus on the inference of a discrete-state Markov process from coarse-grained observations, which is relevant to a number of applied methods [battle2016brokendetailedbalance, lynn2021detailedbalance, nartallokaluarachchi2024decomposing]. In this case, data takes the form of a sequence of coarse-grained states, but bifurcates into two possible situations.
In the first, we have information about $K$ ‘jumps’ and data takes the form, $\{(X_{k},t_{k}):k\in\{1,...,K\}\}$, where $X_{k}$ is the state the process jumped into at ‘jump time’ $t_{k}$. Given this knowledge about the exact jump times, the maximum likelihood estimator (MLE) of the transition intensity from state $j$ to state $i$ is given by,
@@
where $N_{j\rightarrow i}$ is the observed number of jumps from state $j$ to state $i$ that occurred during the $K$ steps, and $T_{j}$ is the holding time i.e. the total time the process spent in state $j$ during the trajectory [Bladt2005statisticalinference].
In the second, more typical, case, data comes in the form $\{X_{t}:t=k\Delta t,k\in\{1,...,K\}\}$, where we observe the process at a series of discrete snapshots, typically equispaced. Given such observations in discrete time, Eq. (64) is no longer the MLE of the transition intensity of the CTMC, but instead an approximation whose accuracy depends on the time-step. Instead, we must infer a discrete-time Markov chain (DTMC). In discrete time, the MLE of the transition probability matrix [^16] A TPM, $P$, defines a DTMC, where the entry $P_{ij}$ is the probability that the process will transition from state $j$ to state $i$ in a single time-step [Ross2019probmodels]. (TPM), $P$, is given by,
@@
where $N_{j\rightarrow k}$ is the number of observed transitions from state $j$ to state kk [^17] In practice, we must also avoid singularities, thus we employ a Bayesian prior. This means we add a ‘pseudo-count’ for all valid transitions [Sahasrabuddhe2025concise].[Bladt2005statisticalinference].
The inference of a CTMC from $P$ becomes identical to the classic embedding problem for Markov matrices [Casanellas2023embedding]. When monitoring a CTMC, with Laplacian $L$, at discrete time-steps, with interval $\Delta t$, the observation is a DTMC with TPM,
@@
Given [^18] In our case, $P=\hat{P}$ is calculated from the observations using Eq. (65). $P$, the embedding problem is to find $L$ such that Eq. (66) is satisfied [Casanellas2023embedding]. Importantly, if $L$ solves the embedding problem for $\hat{P}$, then it is the MLE for the CTMC. However, it need not be unique as the matrix exponential is not injective. One simple condition that guarantees the existence and uniqueness of $L$ is,
@@
A looser condition is that,
@@
where $L$ can be calculated using the unique logarithm of $P$ (see Ref. [Bladt2005statisticalinference] and references therein). The embedding problem has significant implications for applications with coarse temporal observations including in sociology [Singer1976representation], economics [Geweke1986mobility], and evolutionary biology [Verbyla2013embedding], where direct expectation maximisation algorithms are a sensible approach [Bladt2005statisticalinference]. In the case of a diffusion and its coarse-graining, temporal granularity is typically high, thus we opt to use the approximate MLE of Eq. (64), which proved more numerically robust than computing the matrix logarithm.
@@ 5.1. Numerical estimation of the EPR
We sample trajectories from the OU and HO processes (see App. B.1 and B.2 for sampling methods). Fig. 15 shows the stationary distributions of the inferred CTMCs, which are a close approximation of the true density (see Sec. 3). Moreover, we investigate the EPR estimates using trajectories of increasing length and decreasing grid-size. We find that longer trajectories and finer grid-sizes give more accurate estimates of the EPR, which are very stable over 100 iterations. However, in both cases, the inferred EPR is a substantial underestimate of the true EPR, highlighting the challenge of estimating the EPR from finite data, even for fine-grids.
@@ 5.2. Determining if trajectories are from a NESS
A key drawback of inferring the EPR from sampled trajectories is that, by only considering finite data, we may obtain an unreliable estimate of the true EPR. In particular, we may naturally overestimate the irreversibility and flux in the stochastic process, and incorrectly determine that a reversible stationary process is out of equilibrium. For example, if we witness an odd number of transitions between two states, we will naturally observe an asymmetry, even if the true flux between the states is zero. As a result it is necessary to perform ‘surrogate testing’ to determine if the measured irreversibility and EPR is a genuine statistical feature, or an artefact arising from the ‘noise-floor’, the level of irreversibility that exists due to finite data. A typical approach to surrogate testing for time-irreversibility and entropy production is to shuffle trajectories in time, breaking the temporal order, and then re-inferring the EPR to obtain an estimate of the noise-floor [battle2016brokendetailedbalance, lynn2021detailedbalance, nartallokalu2025review]. Whilst this approach works for some direct measures of time-irreversibility, it is unsuitable for this Markov inference approach. This is because during the inference, we only consider transitions between adjacent cells - choosing the resolution of our grid such this accounts for a sufficient number of time-points [^19] For the numerical simulations and data analysis considered in this Section, we typically set the grid-size such that $\approx 70\%$ of time-points are used in the inference.. When the trajectory is randomly shuffled, transitions between neighbouring cells become extremely rare, thus there are not enough valid transitions to infer the surrogate model.

### v2

@@ 4.1.1. The Ornstein-Uhlenbeck process
The Ornstein-Uhlenbeck (OU) process is a multivariate diffusion process that models interacting degrees of freedom under both friction and fluctuations [27]. It is given by the SDE
@@
where $B\in\mathbb{R}^{d}\times\mathbb{R}^{d}$ is the friction matrix and $D=\frac{1}{2}\Sigma\Sigma^{\top}$ is the diffusion matrix. If all eigenvalues of $B$ have positive real-parts, the process converges to a zero-mean Gaussian stationary density given by [^8] As we denote the stationary density as $\pi$, we use $\bm{\pi}$ to denote the mathematical constant.
@@
which corresponds to the multivariate normal distribution $\mathcal{N}_{d}(0,S)$, where $S$ is the covariance matrix satisfying the Lyapunov equation
@@
In general, steady-states of the OU are out of equilibrium. More specifically, the steady-state is in equilibrium if and only if $BD=DB^{\top}$ which implies $S=B^{-1}D$ [38]. Otherwise, the stationary probability flux is given by
@@
where $\mu=DS^{-1}-B$, which then allows for an explicit calculation of the EPR. We first define Q=BSDQ=BS-D, which implies that $\mu=-QS^{-1}$ and $Q=-Q^{\top}$ from the Lyapunov equation, Eq. (43). With Eq. (9) and basic results for Gaussian variables, we can obtain $\Phi=\operatorname{Tr}(B^{\top}D^{-1}Q)$ [27, 12].
@@ 4.1.2. The stochastic Hopf oscillator
The Hopf oscillator [^9] Also known as the Stuart-Landau oscillator [46]. (HO) is a nonlinear process given by the equations
@@
where $\omega$ is the natural frequency and $\sigma$ is the noise intensity. In the absence of noise, for $a<0$ the system decays to the origin which is an attracting fixed point. At $a=0$, the system goes through a Hopf bifurcation, resulting in limit-cycle oscillations for $a>0$ [73]. This system has become a useful model of neural dynamics [54], and hair-cell bundles [69].
To study the NESS of this process in the presence of additive noise, we consider the FP equation in the alternative form
@@
In order to solve for the steady-state, we use that the limit-cycle behaviour of the deterministic system is radially symmetric, and thus assume that the stationary density is also i.e. $\frac{\partial p}{\partial\theta}=\frac{\partial^{2}p}{\partial\theta}=0$, which implies the distribution takes the form
@@
which is solved by a Boltzmann-Gibbs density [^10] It is not immediately obvious why a solution of the form of Eq. (51) solves Eq. (50). However, it is intuitive that the irreversible, oscillatory circulation is driven by the $\theta$ dynamics, which are absent here. This suggests that the dynamics in $r$ may be the negative-gradient of a scalar potential. We can see that $f^{r}=-\frac{dU}{dr}$ for a potential $U=\frac{1}{2}r^{2}(a-\frac{1}{2}r^{2})$. All that remains is to check that $\pi(r)=\exp\left(-2U/\sigma^{2}\right)$ solves Eq. (50). Whilst the stationary density is a Boltzmann-Gibbs density, the steady-state is in nonequilibrium.
@@
where $Z$ is the normalising factor [^11] An alternative approach to solving for the stationary density can be found for the case $a=\omega=1$ in Ref. [79].
@@
Given the stationary density, we can perform the HHD to obtain
@@ 4.1.3. Numerical experiments
Next, by varying the parameter $\theta$, we can vary the EPR, $\Phi=\theta^{2}$, of the underlying OU process. Fig. 5 shows the EPR of the coarse-grained process for a range of $\Delta x$ values as a function of $\theta$. Whilst all the coarse-grained dynamics follow the correct trajectory, increasing the EPR with $\theta$, the smaller step-sizes more closely approximate the true EPR. Moreover, we notice that the EPR of the coarse-grained dynamics may exceed the true EPR. Unlike the EPR of a diffusion observed at a coarse-grained level, the Markov process derived here is not subject to the inequality proved by Esposito in Ref. [13], where it would necessarily lower-bound the true EPR.
@@ 4.2.1. The stochastic van der Pol oscillator
The van der Pol oscillator (VDP) is a nonlinear, non-conservative oscillator that converges to an asymmetric limit-cycle in the absence of noise (see Fig. 9) [73]. We consider a modified, stochastic VDP of the form
@@
For models without an exact stationary density, it is worth additionally verifying that this is the stationary behaviour seen in the long-time-limit, by numerically integrating the FP equation directly. Using the Crank-Nicholson scheme
@@
where $B=I-\frac{\Delta t}{2}L$ and $A=I+\frac{\Delta t}{2}L$ for a time-step $\Delta t$ [39]. We begin with a initial Gaussian density with zero mean and isotropic deviation $\sigma=1$ and integrate up to $T=1.4$ with $\Delta t=0.2$ and $\Delta x=0.05$. As shown in Fig. 10, we see that the FP solution converges to the stationary distribution calculated directly.
@@ 4.2.2. A pair of frustrated Kuramoto oscillators
The Kuramoto oscillator is a model that described the phase of interacting limit-cycle oscillators [37], and is typically used to model synchronisation. It has found a range of applications describing neural dynamics [54], the flashing of fireflies, or the voltage oscillations of Josephson junctions [73].
We first consider the case of a single, isolated Kuramoto oscillator
@@
oscillating at a natural frequency of $\omega>0$ where $\theta(t)\in[0,2\bm{\pi})$. More explicitly, we arrive at the FP equation
@@
Next we focus on a system with a pair of Kuramoto oscillators with additive noise. In addition, their interaction is frustrated, a feature that can lead to nonequilibrium behaviour [28]. We consider
@@
In the first panel of Fig. 14, we calculate the EPR as a function of various values of the coupling strengths. Whilst the system remains out of equilibrium for symmetric coupling, due to the internal oscillations, asymmetry in coupling strength drives the system further from equilibrium, in keeping with previous results on non-reciprocity in collective dynamics [21, 52]. On the other hand, from the second panel, we see that only the absolute value, not the asymmetry, of natural frequencies drives the system from equilibrium, by increasing the speed of the collective oscillation. Finally, in the third panel, we investigate the effect of frustration and noise. Unsurprisingly, noise reduces the EPR by increasing reversible motion. Interestingly, the effect of frustration does not appear to be constant with respect to noise, with the EPR-maximising value of $\alpha$ varying slightly with the noise intensity, $\sigma$.
@@ 5. Statistical inference for discrete-state processes
When analysing NESS in physical and biological systems, we typically do not have access to the function forms of the drift or diffusion. Instead, we must analyse a diffusive process from observations of stochastic trajectories. In order to analyse the observed trajectories of a diffusion, it is common to infer a stochastic model. Whilst some methods attempt to directly infer a SDE [20, 19, 8, 3], we will instead focus on the inference of a discrete-state Markov process from coarse-grained observations, which is relevant to a number of applied methods [2, 42, 53, 56, 36]. In this case, data takes the form of a sequence of coarse-grained states, but bifurcates into two possible situations.
In the first, we have information about $K$ ‘jumps’ and data takes the form, $\{(X_{k},t_{k}):k\in\{1,...,K\}\}$, where $X_{k}$ is the state the process jumped into at ‘jump time’ $t_{k}$. Given this knowledge about the exact jump times, the maximum likelihood estimator (MLE) of the transition rate from state $j$ to state $i$ is given by
@@
where $N_{j\rightarrow i}$ is the observed number of jumps from state $j$ to state $i$ that occurred during the $K$ steps, and $T_{j}$ is the holding time i.e. the total time the process spent in state $j$ during the trajectory [6].
In the second, more typical, case, data comes in the form $\{X_{t}:t=k\Delta t,k\in\{1,...,K\}\}$, where we observe the process at a series of discrete snapshots, typically equispaced. Given such observations in discrete time, Eq. (61) is no longer the MLE of the transition rate of the CTMC, but instead an approximation whose accuracy depends on the time-step. Instead, we must infer a discrete-time Markov chain (DTMC). In discrete time, the MLE of the transition probability matrix [^15] A TPM, $P$, defines a DTMC, where the entry $P_{ij}$ is the probability that the process will transition from state $j$ to state $i$ in a single time-step [60]. (TPM), $P$, is given by
@@
where $N_{j\rightarrow k}$ is the number of observed transitions from state $j$ to state kk [^16] In practice, we must also avoid singularities, thus we employ a Bayesian prior. This means we add a ‘pseudo-count’ for all valid transitions [62].[6].
The inference of a CTMC from $P$ is known as the embedding problem for Markov matrices [11]. When monitoring a CTMC, with Laplacian $L$, at discrete time-steps, with interval $\Delta t$, the observation is a DTMC with TPM
@@
Given $P$, the embedding problem is to find $L$ such that Eq. (63) is satisfied [11]. [^17] In our case, $P=\hat{P}$ is calculated from the observations using Eq. (62). Importantly, if $L$ solves the embedding problem for $\hat{P}$, then it is the MLE for the CTMC. However, it need not be unique as the matrix exponential is not injective. One simple condition that guarantees the existence and uniqueness of $L$ is
@@
A looser condition is that
@@
where $L$ can be calculated using the unique logarithm of $P$ (see Ref. [6] and references therein). The embedding problem has significant implications for applications with coarse temporal observations including in sociology [70], economics [24], and evolutionary biology [77], where direct expectation maximisation algorithms are a sensible approach [6]. In the case of a diffusion and its coarse-graining, temporal granularity is typically high, thus we opt to use the approximate MLE of Eq. (61), which proved more numerically robust than computing the matrix logarithm.
When inferring the transition rates with Eq. (61), we only count transitions between adjacent states, discarding transitions between non-adjacent states. In practice, we choose time and spatial resolution parameters such that at least 70% of observed transitions are valid. In App. E.1, we consider an alternative optimisation-based approach to the embedding problem, which uses all transitions, but show that the results are comparable, whilst the method is significantly less efficient.
@@ 5.1. Numerical estimation of the EPR
We sample trajectories from the OU and HO processes (see App. D.1 and D.2 for sampling methods). Fig. 15 shows the stationary distributions of the inferred CTMCs, which are a close approximation of the true density (see Sec. 3). Moreover, we investigate the EPR estimates using trajectories of increasing length and decreasing grid-size. We find that longer trajectories and finer grid-sizes give more accurate estimates of the EPR, which are very stable over 100 iterations. However, in both cases, the inferred EPR is a substantial underestimate of the true EPR, highlighting the challenge of estimating the EPR from finite data, even for fine-grids.
@@ 5.2. Determining if trajectories are from a NESS
A key drawback of inferring the EPR from sampled trajectories is that, by only considering finite data, we may obtain an unreliable estimate of the true EPR. In particular, we may naturally overestimate the irreversibility and flux in the stochastic process, and incorrectly determine that a reversible stationary process is out of equilibrium. For example, if we witness an odd number of transitions between two states, we will naturally observe an asymmetry, even if the true flux between the states is zero. As a result it is necessary to perform ‘surrogate testing’ to determine if the measured irreversibility and EPR is a genuine statistical feature, or an artefact arising from the ‘noise-floor’, the level of irreversibility that exists due to finite data. A typical approach to surrogate testing for time-irreversibility and entropy production is to shuffle trajectories in time, breaking the temporal order, and then re-inferring the EPR to obtain an estimate of the noise-floor [2, 42, 54]. Whilst this approach works for some direct measures of time-irreversibility, it is unsuitable for this Markov inference approach. This is because during the inference, we only consider transitions between adjacent cells - choosing the resolution of our grid such that this accounts for a sufficient number of time-points. When the trajectory is randomly shuffled, transitions between neighbouring cells become extremely rare, thus there are not enough valid transitions to infer the surrogate model.